According to a Gallup poll, more than half of all Americans drink at least one glass of soda per day. Suppose X is a random variable that represents the number of glasses of soda that a randomly selected American drinks each day. The probability distribution for X is given:

X 0 1 2 3 4
P(X=x) .4 .33 .14 .09 .04

Given a randomly selected American drinks at least two glasses of soda each day what is the probability that they will drink a total of four glasses of soda each day?

Suppose two Americans are selected at random what is the probability that both of them will drink at least two glasses of soda each day?
Answer #1

1)

Given a randomly selected American drinks at least two glasses of soda each day what is the probability that they will drink a total of four glasses of soda each day

=P(X=4|X>=2) =P(X=4)/P(X>=2) =0.04/(0.14+0.09+0.04)=0.1481

2)

probability that both of them will drink at least two glasses of soda each day =P(X>=2)*P(X>=2)=0.27*0.27=0.0729